otal of S hours to devote to ironing shirts and polishing shoes. In past weeks, he has used 3 hours for ironing shirts and 2 hours for polishing shoes. Shoes polished 10 18 24 28 30 Hours per week Shirts ironed 15 28 39 48 Hours per week (a) (2 points) If he chooses to devote a fourth hour to ironing shirts, what is the opportunity cost of doing so? (b) (1 point) If he chooses to devote a fourth hour to ironing shirts, what is the opportunity cost per additional shirt? 2.) The following table shows several points on a small economys production possibilities boundary: Production of Coffee (in tons) Production of Cars (in thousands) 115 95 70 40 Point 6 8 (a) [1 point] Starting from point B, what is the opportunity cost of producing 2,000 more cars? (b) [1 point] Starting from point D, what is the opportunity cost of producing 2,000 more cars?
Answer #1

(1)

(a)

Mortimer can iron 39 shirts in 3 hours and polish 18 shoes in rest 2 hours.

If he devotes 1 more hour to ironing shirts, he will have only (5 - 4) = 1 hour for polishing shoes.

He can iron 48 shirts in 4 hours and polish 10 shoes in rest 1 hour.

Opportunity cost of ironing (48 - 39) = 9 more shirts is (18 - 10) = 8 shoe polish.

(b)

Opportunity cost per additional shirt = (8/9) = 0.89 shoe polish.

NOTE: As per Answering Policy, 1st question is answered.